This question says Solve: 7x^2=168+35x. Thank you so much
Answer by ewatrrr(24785)


Hi,
Solving
7x^2=168+35x.
7x^2-35x-168 = 0
7(x^2-5x-24) = 0
(x -8)(x+3) = 0
(x-8) = 0
x = 8
(x+3)=0
x = -3